6. Use Eulers method to approximate the solution to y = xºy - y? at x = 1.2 when y(0) =1. Use a step size of h=.1.
Answer #1

Doubt in this then comment below..i will help you..

.

please thumbs up for this solution.. thanks..

.

The Eulers method states that Yn+1 = yn + h. f (xn, Yn), where Xn+1 = xn + h. 1 We have that h = „X0 = 0, yo = 1, f(x, y) =Step 6. 1 3 X6 = X5 + h = 1 2 + 10 5 3 y (26) = y 5 = Y6 = 45+hf (35,95) = 0.653779501769385+h --(,10.6 0.653779501769385 85)Step 11. 211 = 210 +h=1+ 1 10 = 11 10 y (211) = y * 911 = 410 + hºf (210, Y10) = 0.584250214360855 +h ·f (1,0.584250214360855